OpenStudy (anonymous):

\[4\sqrt{X^12}\] / \[24\sqrt{X^2}\] its a fraction and in the first sqrt its x^12

OpenStudy (anonymous):

sqrt = power 1/2

OpenStudy (anonymous):

LOL, nearly....

OpenStudy (anonymous):

Use \frac{}{} for a fraction

OpenStudy (anonymous):

4(x^6)/24x= x^6/6x=x^5/6

OpenStudy (anonymous):

is that the answer misterchyme?

OpenStudy (anonymous):

(x^5)/6 yes

OpenStudy (anonymous):

(x^5)/6 is the answer?

OpenStudy (anonymous):

\[\sqrt[4]{x^{12}}=x^3\]

OpenStudy (anonymous):

Though he said sqrt..

OpenStudy (anonymous):

thats the solution for the first one

OpenStudy (anonymous):

Oh, not square root, fourth root.

OpenStudy (anonymous):

So next one is 24th root?

OpenStudy (anonymous):

oh i thought it was 4 times the sqrt

OpenStudy (anonymous):

i mean im supposed to have a simplification for the whole thing

OpenStudy (anonymous):

its a fraction the roots are on top of eachother

OpenStudy (anonymous):

\[\sqrt[24]{x^2}=\sqrt[12]{x}\]

OpenStudy (anonymous):

ok sorry. i guess i am lost.

OpenStudy (anonymous):

thank you guys for the help, its confusing

OpenStudy (anonymous):

ill send a file attached so you can see it

OpenStudy (anonymous):

In English it's the fourth root of x power 12 divided by 24th root of x squared, is that right?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=simplyfy+%284sqrt%28x^12%29%29%2F%2824sqrtx^2%29 this is how i saw it..

OpenStudy (anonymous):

i have no idea estudier

OpenStudy (anonymous):

Look like constants to me

OpenStudy (anonymous):

see i was right! lol

OpenStudy (anonymous):

ohh hahha so whats the answer misterchyme?

OpenStudy (anonymous):

(x^5)/6 final answer

OpenStudy (anonymous):

thanks can you help me with more misterchyme?

OpenStudy (anonymous):

\[\dfrac{4 \sqrt {x^{12}}} {24 \sqrt{x^2}}\]

OpenStudy (anonymous):

Post it as a new question, it will just confuse everyone to put it here..

OpenStudy (deoxna):

\[\sqrt{x^{10}}/6\]

OpenStudy (anonymous):

deox is that the final answer?

OpenStudy (anonymous):

estudier i would if i knew how to attach it to make a new question im new at this

OpenStudy (anonymous):

30x^8sqrt2/5x^2sqrt2 6x^6... i think

OpenStudy (anonymous):

:/ can you help me with this?

OpenStudy (anonymous):

10x^2